OpenStudy (anonymous):

The sum of two numbers is 41 and the difference is 7. What are the numbers?

OpenStudy (anonymous):

17+24=41

OpenStudy (anonymous):

a+b=41 a-b=7 add both eqns 2a=48 a=24 b=17

OpenStudy (anonymous):

i did it with n :) but about the same you did